# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v2

@@ 4.3 Convergence to stationary points
The proof builds upon techniques used in the proof of Theorem 1 in [61]. Let $(\bar{x},\bar{y})$ be an accumulation point of the sequence $\{(x^{k},y^{k})\}$, and let $\{(x^{k_{j}+1},y^{k_{j}+1})\}$ be a convergent subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\rightarrow(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. By Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ Since $\{p_{k}\}$ is bounded, there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$.
@@
$\xi_{y}^{k}$ $\in\nabla_{y}F(x^{k},y^{k+1})+p_{k}\left(\partial_{y}\varphi(x^{k},y^{k+1})-\nabla v_{\gamma}(x^{k},y^{k+1})\right)+\mathcal{N}_{Y}(y^{k+1}),$
@@
$\xi_{x}^{k}$ $:=p_{k}\left(\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-d_{x}^{k}\right)-\frac{p_{k}}{\alpha_{k}}\left(x^{k+1}-x^{k}\right),$ (47)
$\xi_{y}^{k}$ $:=p_{k}\left(\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\right)-\frac{p_{k}}{\beta_{k}}\left(y^{k+1}-\tilde{y}^{k}\right).$
@@
where the last inequality follows from (39). Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and assuming that $\nabla_{x}F,\nabla_{x}f$ and $\nabla_{x}g$ are continuous on $X\times Y$, and $\nabla_{x}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we conclude that $\nabla_{x}\tilde{\psi}_{p_{k}}=\nabla_{x}\tilde{\psi}_{\bar{p}}$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{x}\tilde{\psi}_{p_{k_{j}}}(x^{k_{j}+1},y^{k_{j}+1})-d_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|x^{k_{j}+1}-x^{k_{j}}\|\to 0$ and $\alpha_{k}\geq\underline{\alpha}>0$, we obtain from (47) that $\xi_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
@@
Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and under the assumption that $\nabla_{y}F$ and $\nabla_{y}f$ are continuous on $X\times Y$, and $\nabla_{y}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we have that $\nabla_{y}(\tilde{\psi}_{p_{k}}-g)=\nabla_{y}(\tilde{\psi}_{\bar{p}}-g)$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{y}(\tilde{\psi}_{p_{k_{j}}}-g)(x^{k_{j}},y^{k_{j}+1})-d_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|y^{k_{j}+1}-\tilde{y}^{k_{j}}\|\to 0$ and $\beta_{k}\geq\underline{\beta}>0$, we conclude from (49) that $\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
@@
Finally, from Proposition 3, we know that the limit point $(\bar{x},\bar{y})$ satisfies $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$. Let $\bar{\epsilon}:=\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\in[0,\epsilon]$. Combining this with (50), we have
[…]
@@
As established in Lemma 6, the Ostrowski condition $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$ is satisfied. By combining this result with Theorem 2 and [25, Proposition 8.3.10], we derive the following convergence result:
Corollary 3.
@@
Remark 4.
Our algorithm AGILS offers significant advantages over the MEHA in [42], particularly regarding step-size selection and convergence guarantees. First, AGILS allows a larger, explicit, and easily computable range of step sizes. In contrast, the step sizes for MEHA must be small and lack a clear, practical selection range. This makes AGILS more flexible and easier to tune in practice. Second, AGILS ensures that any accumulation point of the iterates is a feasible KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$. Conversely, MEHA’s convergence analysis only ensures that the stationary residual of penalized problem (12) converges to zero, without guaranteeing the feasibility of the iterates.
@@ 5 Sequential convergence under KL property
This section establishes the sequential convergence of AGILS (Algorithm 1) under the Kurdyka-Łojasiewicz (KL) property. Our analysis focuses on the scenario where the feasibility correction procedure is applied only a finite number of times. We begin by recalling the definition of the KL property. Let $\zeta\in[0,+\infty)$ and $\Phi_{\zeta}$ represent the set of all concave and continuous functions $\phi\ :\ [0,\zeta)\to[0,+\infty)$ that meet the following conditions: (a) $\phi(0)=0$, (b) $\phi$ is $C^{1}$ on $(0,\zeta)$ and continuous at 0, (c) $\phi^{\prime}(s)>0$ for all $s\in(0,\zeta)$.
Definition 2 (Kurdyka-Łojasiewicz property).
@@
Lemma 8 (Uniformized KL property).
Given a compact set $D$ and a proper and lower semicontinuous function $h\ :\ {\mathbb{R}}^{d}\to(-\infty,+\infty]$, suppose that $h$ is constant on $D$ and satisfies the KL property at each point of $D$. Then, there exist $\epsilon$, $\zeta$ and $\phi\in\Phi_{\zeta}$ such that for all $\bar{x}\in D$ and $x\in\left\{x\in{\mathbb{R}}^{d}\,|\,{\rm dist}(x,D)<\epsilon,\ h(\bar{x})<h(x)<h(\bar{x})+\zeta\right\}$,
@@
Assumption 4 is not restrictive. For example, if $g(x,y)=\sum_{i=1}^{p}Q_{i}(x)P_{i}(y)$, where $Q_{i}(x):X\rightarrow\mathbb{R}$ is Lipschitz continuous and strictly positive on $X$, and $P_{i}(y):Y\rightarrow\mathbb{R}$, then $g(x,y)$ satisfies Assumption 4. Specific examples include $g(x,y)=x\|y\|_{1}$ on ++×$\mathbb{R}_{++}\times\mathbb{R}^{m}$ and $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$ on ++$\mathbb{R}^{J}_{++}\times\mathbb{R}^{m}$.
@@
where $\rho_{v_{1}}=\rho_{f_{1}}+\rho_{g_{1}}+c_{\alpha}/4$ and $\tilde{v}_{\gamma}(x,y):={v}_{\gamma}(x,y)+(\rho_{v_{1}}/2)\|x\|^{2}$. By Proposition 2 and under Assumption 3, $\tilde{v}_{\gamma}(x,y)$ is strongly convex with respect to $x$. $H(u,y)$ is the Fenchel conjugate of $\tilde{v}_{\gamma}(x,y)$ with respect to variable $x$ and is differentiable due to the differentiability of ${v}_{\gamma}(x,y)$ and the strong convexity of ${v}_{\gamma}(x,y)$ with respect to $x$. For the sequence $\{(x^{k},y^{k})\}$ generated by AGILS(Algorithm 1), let $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. We have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$ and $\nabla_{y}H(u^{k},y^{k+1})=\nabla_{y}\tilde{v}_{\gamma}(x^{k},y^{k+1})$.
@@
Lemma 9.
Under Assumptions 1, 2, 3 and 4. Let $\epsilon>0$. Let $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ be the sequences generated by AGILS under the same conditions as in Section 4.2, and let $E_{p}(z,u)$ be defined as in (52). In the case where $\tilde{y}^{k}=y^{k}$, there exist constants $a,b>0$ such that
@@
We begin by deriving the first inequality. In this case, $\tilde{y}^{k}=y^{k}$, the inequality (35) simplifies to $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)-\left(1/\beta_{k}-L_{\psi_{y,k}}/2-c_{\beta}/4\right)\left\|y^{k+1}-y^{k}\right\|^{2}+C_{\eta}^{2}/(c_{\beta}\gamma^{2})\zeta_{k}^{2}$, which implies
@@
Next, by Lemma 5, $\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k},\,x^{k+1}-x^{k}\rangle-\left(1/\alpha_{k}-L_{\psi_{x,k}}/2\right)\|x^{k+1}-x^{k}\|^{2}$. Moreover, since the gradients $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$, and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$, $L_{f_{x}}$ and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, we have
@@
By the convexity of $\tilde{v}_{\gamma}$ with respect to $x$ and $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have the identity $-\tilde{v}_{\gamma}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k}\rangle=H(u^{k},y^{k+1})$. Thus, the left-hand side of inequality (58) equals $G_{p_{k}}(z^{k+1})-\langle u^{k},x^{k+1}\rangle+H(u^{k},y^{k+1})$. Furthermore, since $-\tilde{v}_{\gamma}(x^{k},y^{k})\leq-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$, the term $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})$ on the right-hand side of (58) equals $G_{p_{k}}(z^{k})-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$.
@@
[…]
@@
[…]
@@
00 0 $\in\frac{1}{p_{k}}\nabla_{x}F(x^{k},y^{k+1})+\nabla_{x}\varphi(x^{k},y^{k+1})-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})+\frac{x^{k+1}-x^{k}}{\alpha_{k}}+\mathcal{N}_{X}(x^{k+1}),$
00 0 $\in\frac{1}{p_{k}}\nabla_{y}F(z^{k})+\nabla_{y}f(z^{k})+\partial_{y}g(x^{k},y^{k+1})-\frac{\tilde{y}^{k}-\theta^{k}}{\gamma}+\frac{y^{k+1}-\tilde{y}^{k}}{\beta_{k}}+\mathcal{N}_{Y}(y^{k+1}).$
@@
$\tilde{\xi}_{x}^{k}:=$ $\frac{1}{p_{k}}\left(\nabla_{x}F(z^{k+1})-\nabla_{x}F(x^{k},y^{k+1})\right)+\nabla_{x}\varphi(z^{k+1})-\nabla_{x}\varphi(x^{k},y^{k+1})$ (61)
@@
By Assumption 4, for any $\xi_{g}^{k}\in\partial_{y}g(x^{k},y^{k+1})$, there exists $\tilde{\xi}_{g}^{k}\in\mathbb{R}^{m}$ satisfying $\|\tilde{\xi}_{g}^{k}\|\leq L_{g_{y}}\|x^{k+1}-x^{k}\|$ such that ${\xi}_{g}^{k}+\tilde{\xi}_{g}^{k}\in\partial_{y}g(x^{k+1},y^{k+1})$. Consequently, we have $\tilde{\xi}_{y}^{k}\in\partial_{y}G_{p_{k}}(z^{k+1})+\nabla_{y}H(u^{k},y^{k+1})+\mathcal{N}_{Y}(y^{k+1})$, where
$\tilde{\xi}_{y}^{k}:=$ $\frac{1}{p_{k}}\left(\nabla_{y}F(z^{k+1})-\nabla_{y}F(z^{k})\right)+\nabla_{y}f(z^{k+1})-\nabla_{y}f(z^{k})+\tilde{\xi}_{g}^{k}$ (62)
@@
where the last inequality follows from (34) and the $L_{\theta^{*}}$-Lipschitz continuity of $\theta^{*}_{\gamma}(x,y)$ with respect to $y$ (from Lemma 1). For the last component of $\partial E_{p_{k}}(z^{k+1},u^{k})$, note that since $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$. Thus
@@
for some $p_{s}<2$. For example, $s_{k}$ can be chosen as $s_{k}=1/k^{1.1}$. Under these conditions, we can choose a sufficiently large $q>0$ such that $2(1-1/q)>p_{s}$, ensuring that $\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)}<\infty$. Furthermore, by applying the ratio test, as in the proof of Lemma 5, we can show that $\sum_{k=1}^{\infty}k(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})^{2(1-1/q)}<\infty$. Since $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$, it follows that $\sum_{k=1}^{\infty}k\zeta_{k}^{2(1-1/q)}<\infty$.
@@
where $E_{p}(z,u)$ is as defined in (52). Since $w^{q}$ is a KL function, it follows from [58, Theorem 3.6] (although this result considers the generalized concave KL property, it can be directly extended to the KL property using the same proof) that if $E_{p}(z,u)$ is a KL function, then $\tilde{E}_{p}(z,u,w)$ is also a KL function.
Theorem 4.
Under Assumptions 1, 2, 3 and 4, let $\epsilon>0$. Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the same conditions as in Section 4.2, with $s_{k}$ satisfying condition (66). Suppose the sequence $\{(x^{k},y^{k})\}$ is bounded, and for sufficiently large $k$, $p_{k}=\bar{p}$ and the feasibility correction procedure (Algorithm 2) is not applied ( i.e., $\tilde{y}^{k}=y^{k}$). If $E_{\bar{p}}(z,u)$ is a KL function, then the sequence $\left\{x^{k},y^{k}\right\}$ converges to a KKT point of $({\rm VP})_{\gamma}^{\tilde{\epsilon}}$ for some $\tilde{\epsilon}\leq\epsilon$.

### v3

@@ 4.3 Convergence to stationary points
The proof builds upon techniques used in the proof of Theorem 1 in [62]. Let $(\bar{x},\bar{y})$ be an accumulation point of the sequence $\{(x^{k},y^{k})\}$, and let $\{(x^{k_{j}+1},y^{k_{j}+1})\}$ be a convergent subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\rightarrow(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. By Lemma 4.5, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ . Since $\{p_{k}\}$ is bounded, there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$.
@@
$\xi_{y}^{k}$ $\in\nabla_{y}F(x^{k},y^{k+1})+p_{k}\left(\partial_{y}\varphi(x^{k},y^{k+1})-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})\right)+\mathcal{N}_{Y}(y^{k+1}),$
@@
$\xi_{x}^{k}$ $=p_{k}\left(\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-d_{x}^{k}\right)-\frac{p_{k}}{\alpha_{k}}\left(x^{k+1}-x^{k}\right),$ (52)
$\xi_{y}^{k}$ $=p_{k}\left(\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\right)-\frac{p_{k}}{\beta_{k}}\left(y^{k+1}-\tilde{y}^{k}\right).$
@@
where the last inequality follows from (39). Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and assuming that $\nabla_{x}F,\nabla_{x}f$ and $\nabla_{x}g$ are continuous on $X\times Y$, and $\nabla_{x}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2.8), we conclude that $\nabla_{x}\tilde{\psi}_{p_{k}}=\nabla_{x}\tilde{\psi}_{\bar{p}}$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{x}\tilde{\psi}_{p_{k_{j}}}(x^{k_{j}+1},y^{k_{j}+1})-d_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|x^{k_{j}+1}-x^{k_{j}}\|\to 0$ and $\alpha_{k}\geq\underline{\alpha}>0$, we obtain from (52) that $\xi_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
@@
Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and under the assumption that $\nabla_{y}F$ and $\nabla_{y}f$ are continuous on $X\times Y$, and $\nabla_{y}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2.8), we have that $\nabla_{y}(\tilde{\psi}_{p_{k}}-g)=\nabla_{y}(\tilde{\psi}_{\bar{p}}-g)$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},\tilde{y}^{k_{j}})=(\bar{x},\bar{y})$ and $\zeta_{k}\rightarrow 0$, it follows that $\nabla_{y}(\tilde{\psi}_{p_{k_{j}}}-g)(x^{k_{j}},y^{k_{j}+1})-d_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with $\|y^{k_{j}+1}-\tilde{y}^{k_{j}}\|\to 0$ and $\beta_{k}\geq\underline{\beta}>0$, we conclude from (54) that $\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
@@
Finally, from Proposition 4.7, we know that the limit point $(\bar{x},\bar{y})$ satisfies $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$. Let $\bar{\epsilon}:=\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\in[0,\epsilon]$. Combining this with (55), we have
$\left\{\begin{aligned} &0\in\nabla F(\bar{x},\bar{y})+\bar{p}\left(\partial\varphi(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})\right)+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}),\\ &\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\bar{\epsilon}\leq 0,\qquad\bar{p}\left(\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\bar{\epsilon}\right)=0,\end{aligned}\right$. (56)
@@
As established in Lemma 4.5, the Ostrowski condition $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$ is satisfied. By combining this result with Theorem 4.10 and [25, Proposition 8.3.10], we derive the following convergence result:
Corollary 4.11.
@@
Remark 4.12.
Our algorithm AGILS offers significant advantages over the MEHA in [42], particularly regarding step-size selection and convergence guarantees. First, AGILS allows a larger, explicit, and easily computable range of step sizes. In contrast, the step sizes for MEHA must be small and lack a clear, practical selection range. Second, AGILS ensures that any accumulation point of the iterates is a feasible KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$. Conversely, MEHA’s convergence analysis only ensures that the stationary residual of penalized problem (12) converges to zero, without guaranteeing the feasibility of the iterates.
@@ 5 Sequential convergence under KL property
This section establishes the sequential convergence of AGILS (Algorithm 1) under the Kurdyka-Łojasiewicz (KL) property. Our analysis focuses on the scenario where the feasibility correction procedure is applied only a finite number of times. (see Proposition 4.9 for a sufficient condition ensuring this). We begin by recalling the definition of the KL property. Let $\zeta\in[0,+\infty)$ and $\Phi_{\zeta}$ represent the set of all concave and continuous functions $\phi\ :\ [0,\zeta)\to[0,+\infty)$ that meet the following conditions: (a) $\phi(0)=0$, (b) $\phi$ is $C^{1}$ on $(0,\zeta)$ and continuous at 0, (c) $\phi^{\prime}(s)>0$ for all $s\in(0,\zeta)$.
Definition 5.1 (Kurdyka-Łojasiewicz property).
@@
Lemma 5.2 (Uniformized KL property).
[…]
@@
Assumption 5.3 is not restrictive. For example, if $g(x,y)=\sum_{i=1}^{p}Q_{i}(x)P_{i}(y)$, where $Q_{i}(x):X\rightarrow\mathbb{R}$ is Lipschitz continuous and strictly positive on $X$, and $P_{i}(y):Y\rightarrow\mathbb{R}$, then $g(x,y)$ satisfies Assumption 5.3. Specific examples include $g(x,y)=x\|y\|_{1}$ on $\mathbb{R}_{++}\times\mathbb{R}^{m}$ and $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$ on $\mathbb{R}^{J}_{++}\times\mathbb{R}^{m}$.
@@
where $\rho_{v_{1}}=\rho_{f_{1}}+\rho_{g_{1}}+c_{\alpha}/4$ and $\tilde{v}_{\gamma}(x,y):={v}_{\gamma}(x,y)+(\rho_{v_{1}}/2)\|x\|^{2}$. By Proposition 2.8 and under Assumption 2.3, $\tilde{v}_{\gamma}(x,y)$ is strongly convex with respect to $x$. $H(u,y)$ is the Fenchel conjugate of $\tilde{v}_{\gamma}(x,y)$ with respect to variable $x$ and is differentiable due to the differentiability of ${v}_{\gamma}(x,y)$ and the strong convexity of ${v}_{\gamma}(x,y)$ with respect to $x$. For the sequence $\{(x^{k},y^{k})\}$ generated by AGILS (Algorithm 1), let $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. We have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$ and $\nabla_{y}H(u^{k},y^{k+1})=\nabla_{y}\tilde{v}_{\gamma}(x^{k},y^{k+1})$.
@@
Lemma 5.4.
Under Assumptions 2.1, 2.2, 2.3 and 5.3. Let $\epsilon>0$. Let $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ be the sequences generated by AGILS under the same conditions as in Section 4.2, and let $E_{p}(z,u)$ be defined as in (57), In the case where $\tilde{y}^{k}=y^{k}$, there exist constants $a,b>0$ such that
@@
We begin by deriving the first inequality. In this case, $\tilde{y}^{k}=y^{k}$, the inequality (35) simplifies to $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)-\left(1/\beta_{k}-L_{\psi_{y,k}}/2-c_{\beta}/4\right)$
$\left\|y^{k+1}-y^{k}\right\|^{2}+C_{\eta}^{2}/(c_{\beta}\gamma^{2})\zeta_{k}^{2}$, which implies
@@
Next, by Lemma 4.4, $\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k},\,x^{k+1}-x^{k}\rangle-\left(1/\alpha_{k}-L_{\psi_{x,k}}/2\right)\|x^{k+1}-x^{k}\|^{2}$. Moreover, since the gradients $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$, and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$, $L_{f_{x}}$ and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, we have
@@
By the convexity of $\tilde{v}_{\gamma}$ with respect to $x$ and $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have the identity $\tilde{v}_{\gamma}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k}\rangle=H(u^{k},y^{k+1})$. Thus, the left-hand side of inequality (63) equals $G_{p_{k}}(z^{k+1})-\langle u^{k},x^{k+1}\rangle+H(u^{k},y^{k+1})$. Furthermore, since $\tilde{v}_{\gamma}(x^{k},y^{k})\leq-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$, the term $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})$ on the right-hand side of (63) equals $G_{p_{k}}(z^{k})-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$.
@@
For $\alpha_{k}\in[\underline{\alpha},\ 2/(L_{\psi_{x,k}}+c_{\alpha})]$ and recalling $L_{E_{x,k}}=L_{\psi_{x,k}}+c_{\alpha}/4$, we have
@@
Similarly, for $\beta_{k}\in[\underline{\beta},\ 2/(L_{\psi_{y,k}}+c_{\beta})]$, the second coefficient satisfies
@@
$0$ $\in\frac{1}{p_{k}}\nabla_{x}F(x^{k},y^{k+1})+\nabla_{x}\varphi(x^{k},y^{k+1})-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})+\frac{x^{k+1}-x^{k}}{\alpha_{k}}+\mathcal{N}_{X}(x^{k+1}),$
$0$ $\in\frac{1}{p_{k}}\nabla_{y}F(z^{k})+\nabla_{y}f(z^{k})+\partial_{y}g(x^{k},y^{k+1})-\frac{\tilde{y}^{k}-\theta^{k}}{\gamma}+\frac{y^{k+1}-\tilde{y}^{k}}{\beta_{k}}+\mathcal{N}_{Y}(y^{k+1}).$
@@
$\tilde{\xi}_{x}^{k}$= $\frac{1}{p_{k}}\left(\nabla_{x}F(z^{k+1})-\nabla_{x}F(x^{k},y^{k+1})\right)+\nabla_{x}\varphi(z^{k+1})-\nabla_{x}\varphi(x^{k},y^{k+1})$ (66)
@@
By Assumption 5.3, for any $\xi_{g}^{k}\in\partial_{y}g(x^{k},y^{k+1})$, there exists $\tilde{\xi}_{g}^{k}\in\mathbb{R}^{m}$ satisfying $\|\tilde{\xi}_{g}^{k}\|\leq L_{g_{y}}\|x^{k+1}-x^{k}\|$ such that ${\xi}_{g}^{k}+\tilde{\xi}_{g}^{k}\in\partial_{y}g(x^{k+1},y^{k+1})$. Consequently, we have $\tilde{\xi}_{y}^{k}\in\partial_{y}G_{p_{k}}(z^{k+1})+\nabla_{y}H(u^{k},y^{k+1})+\mathcal{N}_{Y}(y^{k+1})$,
$\tilde{\xi}_{y}^{k}$= $\frac{1}{p_{k}}\left(\nabla_{y}F(z^{k+1})-\nabla_{y}F(z^{k})\right)+\nabla_{y}f(z^{k+1})-\nabla_{y}f(z^{k})+\tilde{\xi}_{g}^{k}$ (67)
@@
where the last inequality follows from (34) and the $L_{\theta^{*}}$-Lipschitz continuity of $\theta^{*}_{\gamma}(x,y)$ with respect to $y$ (from Lemma 2.9). For the last component of $\partial E_{p_{k}}(z^{k+1},u^{k})$, note that since $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$. Thus
@@
for some $p_{s}<2$. For example, $s_{k}$ can be chosen as $s_{k}=1/k^{1.1}$. Under these conditions, we can choose a sufficiently large $q>0$ such that $2(1-1/q)>p_{s}$, ensuring that $\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)}<\infty$. Furthermore, by applying the ratio test, as in the proof of Lemma 4.4, we can show that $\sum_{k=1}^{\infty}k((\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0}))^{2(1-1/q)}<\infty$. Since $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$, it follows that $\sum_{k=1}^{\infty}k\zeta_{k}^{2(1-1/q)}<\infty$.
@@
where $E_{p}(z,u)$ is as defined in (57). Since $w^{q}$ is a KL function, it follows from [59, Theorem 3.6] (although this result considers the generalized concave KL property, it can be directly extended to the KL property using the same proof) that if $E_{p}(z,u)$ is a KL function, then $\tilde{E}_{p}(z,u,w)$ is also a KL function.
We now present the global convergence theorem.
Theorem 5.5.
Under Assumptions 2.1, 2.2, 2.3 and 5.3, let $\epsilon>0$. Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the same conditions as in Section 4.2, with $s_{k}$ satisfying condition (71). Suppose the sequence $\{(x^{k},y^{k})\}$ is bounded, and for sufficiently large $k$, $p_{k}=\bar{p}$ and the feasibility correction procedure (Algorithm 2) is not applied ( i.e., $\tilde{y}^{k}=y^{k}$). If $E_{\bar{p}}(z,u)$ is a KL function, then the sequence $\left\{x^{k},y^{k}\right\}$ converges to a KKT point of $({\rm VP})_{\gamma}^{\tilde{\epsilon}}$ for some $\tilde{\epsilon}\leq\epsilon$.